During 2021, its first year of operations, XYZ Company produced 25,000
units and sold 19,000 units. During 2022, XYZ Company produced 30,000
units and sold 32,000 units. The following information was taken from
XYZ's accounting records for 2021 and 2022:

                                               2021            2022
Direct materials cost per unit ............    $18             $17
Direct labor cost per unit ................    $16             $21
Variable overhead cost per unit ...........    $7              $9
Variable selling & admin cost per unit ....    $4              $6
Fixed overhead (total cost) ...............    $105,000        $144,000
Fixed selling & admin (total cost) ........    $88,000         $96,000

Assume the selling price of XYZ Company's product was $64 per unit for
both years.

Calculate XYZ Company's 2022 gross profit using absorption costing.
Assume XYZ Company employs a LIFO inventory cost flow assumption. 
Answer #1

Unit product cost 2021 = (105000/25000)+18+16+7 = 45.20

Unit product cost 2022 = (144000/30000)+17+21+9 = 51.80

Sales (2022) = 32000*64 = 2048000

Cost of goods sold = (30000*51.80+2000*45.20) = 1644400

Gross profit = Sales-Cost of goods sold = 2048000-1644400 = 403600